OpenStudy (anonymous):

if sinx=.42, what is cos x

OpenStudy (ttop0816):

You can make use of the relationship sin^2 + because^2 = 1 to solve for because. You get because^2 = 1 - sin^2 because = Sqrt[1 - sin^2] For your value of sin, this is because = Sqrt[1 - .42^2] because = Sqrt[1 - .1764] because = Sqrt[.8236] because ~~ .9075 The value of because(x) is approximately .9075

OpenStudy (ttop0816):

if you round up the answer it would be 0.908 (:

OpenStudy (anonymous):

easier way is to find \[\sin^{-1} .42 \]

OpenStudy (anonymous):

you should get x=24.835 approx.

OpenStudy (anonymous):

then find cos 24.835

OpenStudy (anonymous):

and like she said it will be approx. .908

OpenStudy (debbieg):

Except I am not sure how you can get the answer better than to within a \(\pm\), without knowing what quadrant you are in....?

OpenStudy (debbieg):

If x is in Q2, then cos(x)<0. But you both have valid ways for finding it (although when I give my students this kind of problem, it is usually when they are NOT permitted a calculator (but I don't give decimal values for the known function value). So I would give something like, if sin(x)=2/5 and x is in Q2, find cos(x). You can't be sure of the sign on cos(x), without knowing the quadrant.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."